[figure: 9.7 cm]

NEW S = 1/2 ALTERNATING CHAIN COMPOUND HIGH PRESSURE FORM OF (VO)_{2}P_{2}O_{7}

M. Azuma^{a}, T. Saito^{a}, Z. Hiroi ^{b} and M. Takano ^{a}

aInstitute for Chemical Research, Kyoto University, Uji, Kyoto 611-0011, Japan bInstitute for Solid State Physics, University of Tokyo, Roppongi 7-22-1, Minato-ku,

Tokyo 106-8666, Japan

(VO)_{2}P_{2}O_{7}, known as a quantum spin chain compound, was found to undergo a structural transition when treated at 2 GPa and 700 ^{o}C. The high-pressure phase as studied by powder X-ray and neutron diffractions comprises a unique kind of S = 1/2 Heisenberg alternating antiferromagnetic chain, in contrast with the ambient pressure phase containing two crystallographically different chains. The magnetic susceptibility and the high-field magnetization data consistently showed the presence of a spin gap of about 26 K as expected from the structure. Single crystal sample was also obtained by slow cooling the melt at 3 GPa.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : M. Azuma

Address for correspondence : Institute for Chemical Research, Kyoto University, Uji, Kyoto 611-0011, Japan

Email address : masaki@scl.kyoto-u.ac.jp